In square GEOM, the coordinates of G are ( 2,-2) In the coordinates of O are (-4,2). Determine and state the coordinates of vertices E and M
1
Expert's answer
2019-05-08T12:58:12-0400


The diagonals of a square bisect each other and meet at 90°. Therefore, the coordinates of crossing-point of the diagonals coincide with the coordinates of bisection point of GO, which can be found via averaging of coordinates of points G and O

and equal (-1,0).

The diagonals of a square are equal, so projection of GO to X axis must be equal to projection of EM to Y axis and vice versa.

Hence, the coordinates of E are (-1 - "\frac{GO_y}{2}" ,0 - "\frac{GO_x}{2}" ),

where "GO_y" - the projection of GO to Y axis, that equals 4,

"GO_x" - the projection of GO to X axis, that equals 6.

So, coordinates of E are (-3,-3).


The coordinates of M are (-1 + "\frac{GO_y}{2}" , 0 + "\frac{GO_x}{2}").

So, coordinates of M are (1,3).